OpenStudy (anonymous):

HELPP graph y=4/3cos3/2x

OpenStudy (anonymous):

The graph should look like the file I attached.

OpenStudy (anonymous):

make sure that your equation is exactly like the following:\[\frac{4}{ 3 }\cos(\frac{ 3 }{ 2x })\]

OpenStudy (anonymous):